### v2

# MMAC-Copilot: Multi-modal Agent Collaboration Operating System Copilot

Abstract
Autonomous virtual agents are often limited by their singular mode of interaction with real-world environments, restricting their versatility. To address this, we propose the Multi-Modal Agent Collaboration framework (MMAC-Copilot), a framework utilizes the collective expertise of diverse agents to enhance interaction ability with operating systems. The framework introduces a team collaboration chain, enabling each participating agent to contribute insights based on their specific domain knowledge, effectively reducing the hallucination associated with knowledge domain gaps. To evaluate the performance of MMAC-Copilot, we conducted experiments using both the GAIA benchmark and our newly introduced Visual Interaction Benchmark (VIBench). VIBench focuses on non-API-interactable applications across various domains, including 3D gaming, recreation, and office scenarios. MMAC-Copilot achieved exceptional performance on GAIA, with an average improvement of 6.8% over existing leading systems. Furthermore, it demonstrated remarkable capability on VIBench, particularly in managing various methods of interaction within systems and applications. These results underscore MMAC-Copilot’s potential in advancing the field of autonomous virtual agents through its innovative approach to agent collaboration.
[…]
1888
1 Introduction
Figure 1: Illustrates the sequence of actions undertaken by MMAC-Copilot in response to the user’s request "Play <Love Story> on Spotify". (A) Represents the initial state where the "Planner" formulates preliminary plans. . (B) "Programmer" launches the application using a bash command. (C) "Viewer" click the search box and enter <Love Story>. (D) "Viewer" click the play button.
Large Language Models (LLMs) [1, 2, 3, 12, 27, 37, 38] has demonstrated unprecedented capabilities to tackle complex tasks that require human-like reasoning, decision-making, and collaborative efforts [11, 16, 28, 29, 45]. Soon afterwards, Large Vision Models (LVMs) [7, 21, 23, 24, 25, 54] expand beyond traditional text-based processing by integrating a visual dimension into LLMs. The applications of LVMs have reached diverse domains, with one notable example being their pivotal usage in autonomous virtual agents [44].
Usually, autonomous virtual agents require a unified interaction interface and a high degree of generalization to adapt to the challenges posed by the numerous different applications within an operating system. Although previous works [35, 44, 51] that utilize scripts and the planning abilities of LLMs have demonstrated success in some applications there are still some challenges: First, the acquisition of real-world environmental information relies on a single text modality which lead to a limited range of applications. Second, the hallucinations caused by the knowledge domain gap could lead to incorrect planning by virtual agents, resulting in extremely severe consequences.
To engage with the initial challenge, we introduce MMAC-Copilot, a framework designed to promote information exchange and collaboration among multi-modal agents. Specifically, we created five key agents: Planner, an expert in making plans, strategic thinking and resource allocation,ensuring that the collaborative efforts of all agents are directed to achieve the target; Librarian, proficient in managing question-answering tasks and API in retrieving information; Programmer, specializing in coding tasks, particularly skilled at executing Bash commands and Python programs; Viewer, focused on understanding image content for visual information and execute click interaction; and Video Analyst, dedicated to analyzing video content and extracting key events. In addition, there is an auxiliary agents: the Mentor, supervising system operations and agent interactions to ensure smooth functionality and integrity. When receiving a user query, Planner formulates an initial plan based on the user query’s requirements. It then engages with other agents through a team collaboration chain, allowing for a dynamic exchange of information and insight. As the plan progresses, it is continuously adapted based on the insights of each specialized agent.
In resolving the second dilemma, we introduce the team collaboration chain, allowing participating agents to adapt the initial plan crafted based on their domain expertise. Refer to the scenario depicted in Figure 3, where the user request is "Open the Discord application and send Hi to Dylan Li". Planner firstly delineates five sub-tasks: "Open Discord", "Navigate to friends", "Search for Dylan Li", "Open Chat with Dylan Li", and "Send a Message", all of which are assigned to Programmer. Upon completion of the first subtask by the Programmer, Mentor observes that Discord has been successfully launched and notes that "Dylan is already present in the application." Planner, upon receiving this feedback, revises the initial plan to "Open chat with Dylan" and "Send the message", and then reassigns it to Programmer. In this instance, Planner overlooks the fact that programmer needs to specifically integrate Discord’s proprietary API to accomplish these tasks. Adapting to the proprietary interfaces of thousands of applications to achieve universality is clearly impractical. Furthermore, unforeseen scenarios such as pop-up notifications or advertisements may disrupt task execution. Programmer fails to execute and returns the message, "Sorry, I don’t know how to do this." Upon receiving this message, Planner reallocates the task to Viewer. Viewer decomposes the coarse-grained task into finer-grained tasks: "click on Dylan", "type Hi", and "return the message", which are then successfully executed. Additionally, each subtask completion is reviewed by Mentor, and upon approval, the process advances to the next round. The process concludes after all rounds are completed.
We assess MMAC-Copilot using two benchmarks: GAIA [26] and our newly introduced Visual Interaction Benchmark (VIBench). GAIA is designed to test the abilities of general artificial intelligence assistants. The MMAC-Copilot performed in level-1, level-2, and level-3 tasks with scores of 45.16, 20.75, and 6.12, respectively. It outperforms previous methods by 6.8%. VIBench, specifically developed for this study, tests the framework’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office environments. The results exhibits outstanding adaptability, handling interfaces with a high degree of accuracy, enhancing interaction within digitally environments.
In summary, our main contributions are as follow:
  • We propose MMAC-Copilot, a collaborative framework involving specialized agents designed to navigate and perform general operating system tasks. MMAC-Copilot enhances the interaction capabilities of autonomous virtual agents with operating systems by leveraging multi-modality in processing to tasks. The framework consists of distinct modal agents: "Librarian" for question-answering and information retrieval, "Programmer" for executing code and bash commands, "Viewer" for interpreting visual information, and "Video Analyst" for analyzing video content. By leveraging the collaboration of modal agents, MMAC-copilot can handle a diverse range of tasks with enhanced accuracy.
  • To mitigate the incorrect planning issue caused by LLMs’ hallucination, we introduce the team collaboration chain which allows participating agents to adjust the initial plan based on their domain expertise into MMAC-Copilot. It has demonstrated adaptability in handling a variety of case.
  • We develop the Visual Interaction Benchmark (VIBench), specifically designed to assess the system’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office scenarios. VIBench complements existing benchmarks by focusing on the agents’ ability to interact with dynamic and visually complex interfaces, thus providing a more strict test of their practical interaction capabilities.
2 Related work
In this section, we review some related works, with a focus on the domains of LLM-based agents and Autonomous virtual agents.
Figure 2: An overview of MMAC-Copilot, below is the concrete definition of each agent.
2.1 LLM-based agents
The sunrise field of LLM-based agents, highlighted by projects such as AutoGPT [33] and TaskWeaver [30], marks a significant milestone in the field of LLM research. These agents adopt a human-like approach to problem-solving by decomposing complex tasks into manageable subtasks. The capability demonstrates a profound advancement in planning, observation, and responsive actions, closely mirroring human cognitive processes [38, 40, 45].
The LangChain Agent [36] further extends these capabilities by enabling LLMs to sequence actions, thus expanding the potential for automation across various domains such as customer service, logistics, and content creation. This strategic utilization of LLMs underscores a move towards more sophisticated, context-aware AI systems.
Developments in multi-agent systems, such as AutoGen [43], MetaGPT [17], and AutoAgents [8], introduce a novel dimension to the field by enabling collaboration and competition among agents. This architecture leverages the unique strengths of individual agents, enhancing system-wide efficiency and adaptability in complex scenarios. Such collaborative frameworks are proving essential for large-scale applications in sectors like healthcare [10, 20], finance [14], and urban planning [4].
The versatility and efficacy of these LLM-based agents are highlighted in a range of application domains: In robotics, agents assist in both navigational tasks and complex manipulations, contributing to more autonomous and adaptive robotic systems [12]. Web manipulation and data extraction have seen significant enhancements, enabling better data-driven decision-making and user interaction models [49, 55]. The gaming industry benefits from more dynamic AI opponents and game management systems, which improve user engagement and game design [13, 35]. Automated data analysis tools powered by LLM agents offer sophisticated insights and predictions, facilitating more effective data handling in business and research [39, 53].
2.2 Autonomous virtual agents
The application of LVM systems for interfacing with Graphical User Interfaces (GUIs) in digital applications have been a key area of research. Wonderland [46] harnessed the capabilities of GPT-4V which delineated by Yang et al. [47] to navigate through mobile application interfaces by analyzing screenshots. Concurrently, the AppAgent [48] utilizes GPT-4V to simulate the actions of mobile users, facilitation the execution of tasks in mobile applications through the analysis of device snapshots. Moreover, the MobileAgent incorporates Optical Character Recognition (OCR) technology to enhance the capabilities of of GPT-4V within a similar mobile agent architecture. This strategic enhancement allows MobileAgent to attain task completion rates on par with human operators. In contrast, CogAgent [18] constructs a specifically engineered LVM for interpreting and navigating GUIs. UFO [51] utilizes the windows inspect tool to acquire component and control information from system applications, providing it as context to GPT-4V to fully leverage its planning capabilities.
3 MMAC-Copilot Framework
The workflow of MMAC-Copilot is designed to leverage the specialized capabilities of each agent in a sequential and collaborative manner. Here, we present an algorithmic overview in Algorithm 1. In section 3.1, we discuss the specialization of agents within the framework, each equipped with distinct roles and tools. In section 3.2, we delve into communications protocols which bolster accuracy in agents collaborations. In section 3.3, we introduce the team collaboration chain, dynamic framework that utilizes continuous feedback and iterative improvements to adapt framework strategies in real-time.
Algorithm 1 Workflow of MMAC-Copilot
1:User request $R$
2:Task completion that satisfies $R$
3:Initialize system state $S_{0}$
4:Planner formulates initial plan $P_{0}$ based on $R$
5:for each subtask $s$ in $P_{0}$ do
6:     Assign $s$ to appropriate agent $A$
7:     $A$ executes $s$, updates system state to $S_{i}$
8:     Mentor checks $S_{i}$; provides feedback $F_{i}$
9:     if $F_{i}$ indicates adjustment needed then
10:         Planner revises $P_{i}$ based on $F_{i}$
11:         Go to step 4 with revised $P_{i}$
12:     end if
13:end for
14:Verify if $S_{i}$ satisfies $R$
15:if not satisfied then
16:     Retry
17:else
18:     Break loop and finalize task
19:end if
3.1 Specialization of agents
Empirical studies suggest that diversity within human groups contributes to a variety of perspectives, which in turn enhances the group’s performance across various tasks [6, 41, 42]. Also, contemporary research indicates that assigning specific roles to autonomous agents can improve their effectiveness [22, 29, 31]. Thus, in our framework, we have defined six agents, namely: Planner, Librarian, Programmer, Viewer, Video Analyst, and Mentor, as shown in Figure 2. The followings are detailed introductions to each agent:
  • Planner: Employs GPT-4 to strategically manage and allocate tasks among agents, optimizing workflow efficiency. In contrast to other frameworks, our Planner is not required to decompose tasks in to indivisible atomic subtasks. We recognize Planner operating solely in the textual modality may not effectively integrate visual information such as UI elements, potentially leading to hallucination. Therefore, when Planner identifies a subtask which necessitates the incorporation of visual information, it provides only a coarse outline of the subtask. It allows each agent to leverage its own expertise, enhancing the adaptability of the overall framework.
  • Librarian: Utilizes GPT-4 and API for information retrieval, enabling it to answer queries and provide foundational knowledge. It possesses the capability to rapidly process vast amounts of information, thereby supporting user requests and providing related context with other agent.
  • Programmer: Responsible for coding and executing scripts, including Python and Bash commands, to interact with software environments directly. The Programmer’s role is crucial in implementing precise operating systems interactions that adapt to diverse requirements. Key to the Programmer’s efficacy is its ability to refine existing code through a mechanism designed to diagnose and resolve issues effectively. When code errors are identified, Programmer analyzes the source of failure, applies necessary modifications, and maintains the original code structure as much as possible. This refined process includes a detailed explanation of identified issues, enhancing the clarity and functionality of the code without introducing new exception, and improving the handling of existing ones. Furthermore, Programmer can generate accurate Python invocation statements. By deeply understanding the context of the provided Python class, analyzing task descriptions, and interpreting method parameter details to ensure the correct application. The result is a syntactically correct Python statement that calls the methods with appropriate parameters. Lastly, Programmer’s capabilities are complemented by a critical evaluation mechanism. This mechanism assesses the refined and executed code against the user’s specified task requirements. It evaluates the code’s functionality, alignment with the task, and feedback provided by the environment, culminating in a comprehensive reasoning process. The interactions within the Programmer’s mechanism can be mathematically represented as follows:
[…]
    where:
      $C_{\text{init}}$ is the initial code.
      $E$ represents the error analysis and feedback.
      $R$ is refinement mechanism.
      $C_{\text{mod}}$ is the modified code.
      $\delta$ includes environmental variables such as the current working directory or system configurations.
      $X$ represents execution function.
      Out includes the runtime output and any runtime errors.
      $T$ is the original task requirements.
      $V$ represents the evaluation mechanism.
      Judge is a boolean indicating if the task is accomplished.
      Score is an integer rating the code’s generality and robustness. If the score is above the specified threshold, the code for this run is retained.
  • Viewer: Integrates GPT-4V for interpreting complex visual information from screenshots. Coupled with SeeClick [9] model, Viewer could translates visual analysis into actionable commands, such as click on identified UI elements, bridging the gap between visual understanding and operating systems interaction.
  • Video Analyst: Uses Gemini Vision to process and analyze video content, extracting critical visual data. It is significant for subtasks which depend on video-based information, enhancing the model’s ability to generalize across different modal. Video Analyst interprets queries, retrieves relevant details from video streams, and delivers accurate responses. These capabilities are crucial for providing contextual insights, assisting other agents in understanding the current statement. In summary, Video Analyst is adept at handling complex scenarios where video data is key.
  • Mentor: Powered by GPT-4V, Mentor provides strategic oversight and troubleshooting support after the execution processes. To be specific, Mentor firstly analyzes screenshots of the current application window to provide detailed description of visible state. And then Mentor identified whether the intended change or commands have successfully impacted the application as expected. Additionally, Mentor generates insights which are derived from a deep analysis of the outcomes and are structured as feedback to be incorporated into the next round.
3.2 Communication Protocols
Following established multi-agent frameworks [17, 22, 28, 52, 56], MMAC-Copilot adopts Structured Communication Interfaces, avoiding the traps of unconstrained nature language communication. Each agent communicates through a defined schema appropriate for their role.
In specific, we define the output format of the agents as JSON, and stipulate the keys which should be output according to the different roles of each agent. Meanwhile, each agent only receives the content of the required key. For example, "observation" key is exclusive output by Mentor and Viewer, as it is specifically to their role in emphasizing visual observations of screenshots. Apart from themselves, only the Planner receives the content of "observation" key, both when formulating initial plans and when revising them. These protocols are designed to prevent common issues such as data ambiguity in complex multi-agent interactions. By standardizing the communication format and defining clear protocols for message acknowledgment, MMAC-Copilot minimizes the risks of miscommunication and ensures which tasks are executed reliably. Additionally, the use of JSON allows for easy integration of new agents as the system scales, supporting the framework’s adaptability and expansion.
Figure 3: Example of MMAC-Copilot. User query is "Open the Discord and Send a message to Dylan Li". Red represents the agent that will take the next action in the messaging process. Brown indicates the current agent performing the task. Green color represents state feedback.
3.3 Team Collaboration Chain
Team collaboration chains leverage the strengths of dispersed management structures to facilitate enhanced cooperation among team members across different domains. By distributing responsibilities and decision-making power, these chains enable a more agile and responsive organizational environment. Such chain has been successfully implemented across multiple domains [5, 19, 32, 50].
In the MMAC-Copilot framework, team collaboration chain is essential for managing difficult tasks. This section delves into the collaborative mechanisms that enable dynamic planning and the iterative refinement of tasks, ensure effective execution and continuous improvement. A concrete demonstration of the collaboration chain in operation can be seen in Figure 3.
3.3.1 Dynamic Planning and Subtask Refinement
During the planning phase, Planner uses input from the Viewer to formulate an initial plan. Initial plan often outline the subtasks in a coarse manner, addressing the dynamic nature of visual environments and the inherent limitations of planning based on textual information alone. This initial planning stage sets the foundation for task execution but requires further refinement to address the specifics of the visual statement.
As the system transitions to the execution phase, agents equipped with specialized modalities, such as the Viewer and Video Analyst, assume pivotal roles. These agents reassess the current visual context and refine the coarsely defined subtasks into detailed, atomic subtasks. This step is crucial for adapting to real-time changes in the visual interface, ensuring that each task is executed with precision based on the latest visual information. This process showcases the strength of team collaboration, where different agents contribute uniquely to the refinement and execution of tasks.
3.3.2 Continuous Feedback and Iterative Improvement
Mentor plays a fundamental role in the team collaboration chain by establishing a continuous feedback loop with the Planner. After the execution of subtasks, Mentor evaluates the current state of the system through detailed analysis of screenshots and the visible state of the application. This evaluation determines whether the intended changes or commands have successfully impacted the application as expected.
If discrepancies or incomplete tasks are identified, Mentor generates insights and structured feedback based on a deep analysis of the outcomes. This feedback is crucial for revising the current plans and is incorporated into the next round of planning. This feedback loop between the Mentor and the Planner not only ensures high-quality task completion but also facilitates adaptive learning, allowing the MMAC-Copilot system to evolve its strategies based on real-world interactions and outcomes.
4 Experiments
Figure 4: Task Completion under VIBench with MMAC-Copilot. Displayed is a sequential illustration of task completion across three distinct categories: (1) interactive gameplay performing skill in Genshin Impact; (2) operate within office application VooV Meeting; (3) recreational operation, exemplified by navigating in Netflix.
4.1 Benchmark and setting
We evaluate the performance of MMAC-Copilot using two distinct benchmarks designed to challenge and assess the capabilities of AI assistants in diverse interaction environments. The first is the General AI Assistant Benchmark (GAIA) [26], which includes 466 rigorous question-answering tasks. GAIA primarily tests AI assistants in a question-answering (QA) format, focusing on their ability to utilize API calls for information retrieval and task execution. However, GAIA restricts the evaluation to textual interactions and API-based actions, which do not fully capture the broader range of GUI interactions that are prevalent in real-world applications.
To address these limitations and to evaluate the agents’ capabilities in a more complex and visually oriented environment, we introduced the Visual Interaction Benchmark (VIBench). VIBench comprises cases from three categories of applications: 3D gaming, recreation, and office environments. Unlike GAIA, VIBench emphasizes the necessity for agents to interact with applications through graphical user interfaces (GUIs) that cannot typically be controlled via standard APIs such as Win32 APIs.
Mathematical Representation of VIBench
The core interaction process in VIBench can be mathematically described as follows: Let $S_{0}$ represent the initial system state which includes the current GUI appearance and available actions. Given a user request $R$, the system must transition through a series of states $S_{1},S_{2},\ldots,S_{n}$ via actions $a_{1},a_{2},\ldots,a_{n}$, ultimately achieving the final state $S_{f}$ that satisfies $R$. This process is represented by the sequence:
$S_{0}\xrightarrow{a_{1}}S_{1}\xrightarrow{a_{2}}S_{2}\xrightarrow{\cdots}S_{n}\xrightarrow{a_{n}}S_{f}$
The primary goal of VIBench is not merely to evaluate the agent’s ability to execute each step efficiently but to assess whether the final state $S_{f}$ satisfies the initial user request $R$. This objective emphasizes the effectiveness of the final outcome over the process, reflecting the practical focus of real-world applications:
Objective:Verify that $\text{Objective}:\text{Verify that }S_{f}\text{ meets }R$
VIBench thus complements GAIA by adding a layer of complexity and realism, challenging agents to not only respond correctly in a QA format but also to effectively navigate and interact within software environments that mirror everyday user challenges.Examples are shown in Figure 4
Setting
To ensure a fair comparison with previous work, we employed the same versions of the ChatGPT API used in prior studies. Specifically, we utilized GPT-4-turbo-1106 and gpt-4-vision-preview. For the Video Analyst, we used gemini-1.5-pro-vision.
4.2 Evaluation Method
For GAIA, each question calls for a specific response: string, number, or floats. Evaluation is conducted through a exact match between the model’s answer and ground truth.
For our VIBench, we adopt human evaluation which follow the SIMA evaluation method [34], recognizing that in numerous instances, task success cannot be automatically inferred. We engage human judges who possess expert-level familiarity with the specific applications, defined as having at least 20 hours of application usage. These experts are asked to assess recorded videos to determine whether the user requests have been accurately executed. During the assessment, human experts are instructed to disregard any irrelevant actions performed by the agents. Their judgments are solely based on whether the stated objectives of the tasks are achieved, without considering the intermediate steps or unrelated actions. Meanwhile, to minimize human subjectivity, we have imposed a maximum limit of 30 execution rounds for task. If the task is not completed within 30 rounds, it is judged as a failure.
4.3 Baseline
For GAIA, we present the performance of GPT-3.5 and GPT-4 with and without manually set plugins, along with results from AutoGPT, Chamonmile and FRIDAY which all use GPT-4 as their backend. The selection of Plugins for GPT-4 is conducted manually, with individual choosing based on specific user query. For comparative purposes, we also include data on human performance sourced from GAIA [26].
For VIBench, we report the performance of UFO and FRIDAY. Notably, since FRIDAY does not natively support the Windows platform, we have modified FRIDAY’s code to adapt it to Windows. Also we set UFO’s Vision model same as our framework.
4.4 Results
The results of our experiments with MMAC-Copilot provide a comprehensive assessment of its capabilities in handling complex, multi-modal tasks with in digital environments. In this session, we present and discuss the findings from our evaluations on both the GAIA benchmark and our newly developed Visual Interaction Benchmark(VIBench).
Table 1: Performance comparison on GAIA. All performance are reported on the private test set.
Model Level 1 Level 2 Level 3 Average
Human* 93.90 91.80 87.30 91.00
GPT-3.5 [1] 4.3 1.89 2.08 2.67
GPT-4 [1] 9.68 1.89 0.00 4.00
AutoGPT4 [33] 15.05 0.63 0.00 5.00
GPT4 Turbo [1] 9.68 6.92 0.00 6.67
GPT-4 Plugins [1] 30.30 9.70 0.00 14.60
Chamomile [15] 16.13 17.61 2.08 14.67
FRIDAY [44] 40.86 20.13 6.12 24.25
Ours 45.16 20.75 6.12 25.91 (+6.8%)
MMAC-Copilot shows significant performance improvement across all levels of GAIA. As illustrated in Table 1, MMAC-Copilot achieved the highest scores in Level 1 (45.16) and Level 2 (20.75) tasks, matching the highest score in Level 3 tasks (6.12). Also on average, it outperformed the closest competing system, FRIDAY, by 6.8% in overall task performance. Given that tasks in GAIA encompass various complex AI assistant capabilities, from basic information retrieval to interactive task requiring multiple steps and integration of external data sources.
Table 2: Evaluation results on VIBench.
Model 3D Gaming Recreation Office Average
UFO [15] 0.00 28.57 15.38 14.65
FRIDAY [44] 31.58 42.86 30.77 35.07
Ours 63.16 69.23 78.57 70.32 (+35.25)
In VIBench as shown in Table LABEL:tab:VIBench, MMAC-Copilot outperformed other systems across all application categories (3D gaming, recreation, and office scenarios). Our model achieved average scores fo 63.16% in 3D gaming, 69.23% in recreation, and 78.57% in office applications, with an overall average of 70.32%. It indicates MMAC-Copilot’s robust capability in navigating and interacting within complex GUI environments, where direct API interaction is not possible. MMAC-Copilot’s superior performance can be attributed to its effective utilization of image and video analysis, enabling precise navigation and interaction through the GUIs of different software.
5 Conclusion
This paper presents MMAC-Copilot, a novel framework designed to enhance the interaction capabilities of autonomous virtual agents with operating systems. Our approach capitalizes of autonomous virtual agents with operating systems. Our approach capitalizes on a team collaboration chain that harnesses the collective expertise of each agents, including "Planner", "Librarian", "Programmer", "Viewer", "Video Analyst", and "Mentor". Each agent brings a unique set of skills and perspectives, enabling an adaptive interaction with complex digital environments. Also MMAC-Copilot innovatively mitigates the issues with hallucination among the agents. By facilitating dynamic and team collaboration among the agents, our framework allows for continuous adaptation and refinement of plans based on environmental changes. The performance of MMAC-Copilot has been tested on GAIA benchmark and our newly introduced VIBench, showing significant improvements over existing state-of-the-art systems.
6 Discussion
We acknowledge that while MMAC-Copilot has achieved a level of generality in application through the integration of GUI models for visual detection, it is still limited by the GUI model’s ability to understand UI components. Consequently, MMAC-Copilot struggles with interpreting complex UI interfaces and performing multi-step operations that depend on such complexity. Additionally, MMAC-Copilot’s capabilities in real-time 3D gaming environments reveal gaps in the spatial understanding, which is cruical for tasks such as navigation, aiming and shooting in 3D environments. The framework’s longer inference time impair the ability to perform tasks with higher temporal sensitivity.
To solve these, we suggest integrating external knowledge databases via online search engines, providing additional context for decision-making processes. This approach will not only help in understanding complex UI layouts but also in recognizing user intentions and environmental cues in real-time games.

### v3

@@
-MMAC-Copilot: Multi-modal Agent Collaboration Operating System Copilot
+MMAC-Copilot: Multi-modal Agent Collaboration Copilot
 Abstract
-Autonomous virtual agents are often limited by their singular mode of interaction with real-world environments, restricting their versatility.
-To address this, we propose the Multi-Modal Agent Collaboration framework (MMAC-Copilot), a framework utilizes the collective expertise of diverse agents to enhance interaction ability with operating systems.
+Large language model agents that interact with PC applications often face limitations due to their singular mode of interaction with real-world environments, leading to restricted versatility and frequent hallucinations.
+To address this, we propose the Multi-Modal Agent Collaboration framework (MMAC-Copilot), a framework utilizes the collective expertise of diverse agents to enhance interaction ability with application.
 The framework introduces a team collaboration chain, enabling each participating agent to contribute insights based on their specific domain knowledge, effectively reducing the hallucination associated with knowledge domain gaps.
-To evaluate the performance of MMAC-Copilot, we conducted experiments using both the GAIA benchmark and our newly introduced Visual Interaction Benchmark (VIBench).
+We evaluate MMAC-Copilot using the GAIA benchmark and our newly introduced Visual Interaction Benchmark (VIBench).
+MMAC-Copilot achieved exceptional performance on GAIA, with an average improvement of 6.8% over existing leading systems.
 VIBench focuses on non-API-interactable applications across various domains, including 3D gaming, recreation, and office scenarios.
-MMAC-Copilot achieved exceptional performance on GAIA, with an average improvement of 6.8% over existing leading systems.
-Furthermore, it demonstrated remarkable capability on VIBench, particularly in managing various methods of interaction within systems and applications.
-These results underscore MMAC-Copilot’s potential in advancing the field of autonomous virtual agents through its innovative approach to agent collaboration.
-[…]
-1888
+It also demonstrated remarkable capability on VIBench.
+We hope this work can inspire in this field and provide a more comprehensive assessment of Autonomous agents.
+The anonymous Github is available at Anonymous Github
+MMAC-Copilot: Multi-modal Agent Collaboration Copilot
+Zirui Song1,2∗Yaohang Li2thanks: Equal contributions.Meng Fang4∗Yanda Li2Zhenhao Chen1 Yuan Huang3Zhecheng Shi3Xiuying Chen1Ling Chen2 1 Mohamed bin Zayed University of Artificial Intelligence 2 University of Technology Sydney 3 Northeastern University 4 University of Liverpool
 1 Introduction
-Figure 1: Illustrates the sequence of actions undertaken by MMAC-Copilot in response to the user’s request "Play <Love Story> on Spotify".
-(A) Represents the initial state where the "Planner" formulates preliminary plans.
-.
-(B) "Programmer" launches the application using a bash command.
-(C) "Viewer" click the search box and enter <Love Story>.
-(D) "Viewer" click the play button.
-Large Language Models (LLMs) [1, 2, 3, 12, 27, 37, 38] has demonstrated unprecedented capabilities to tackle complex tasks that require human-like reasoning, decision-making, and collaborative efforts [11, 16, 28, 29, 45].
-Soon afterwards, Large Vision Models (LVMs) [7, 21, 23, 24, 25, 54] expand beyond traditional text-based processing by integrating a visual dimension into LLMs.
-The applications of LVMs have reached diverse domains, with one notable example being their pivotal usage in autonomous virtual agents [44].
-Usually, autonomous virtual agents require a unified interaction interface and a high degree of generalization to adapt to the challenges posed by the numerous different applications within an operating system.
-Although previous works [35, 44, 51] that utilize scripts and the planning abilities of LLMs have demonstrated success in some applications there are still some challenges: First, the acquisition of real-world environmental information relies on a single text modality which lead to a limited range of applications.
-Second, the hallucinations caused by the knowledge domain gap could lead to incorrect planning by virtual agents, resulting in extremely severe consequences.
-To engage with the initial challenge, we introduce MMAC-Copilot, a framework designed to promote information exchange and collaboration among multi-modal agents.
-Specifically, we created five key agents: Planner, an expert in making plans, strategic thinking and resource allocation,ensuring that the collaborative efforts of all agents are directed to achieve the target; Librarian, proficient in managing question-answering tasks and API in retrieving information; Programmer, specializing in coding tasks, particularly skilled at executing Bash commands and Python programs; Viewer, focused on understanding image content for visual information and execute click interaction; and Video Analyst, dedicated to analyzing video content and extracting key events.
-In addition, there is an auxiliary agents: the Mentor, supervising system operations and agent interactions to ensure smooth functionality and integrity.
-When receiving a user query, Planner formulates an initial plan based on the user query’s requirements.
-It then engages with other agents through a team collaboration chain, allowing for a dynamic exchange of information and insight.
-As the plan progresses, it is continuously adapted based on the insights of each specialized agent.
-In resolving the second dilemma, we introduce the team collaboration chain, allowing participating agents to adapt the initial plan crafted based on their domain expertise.
-Refer to the scenario depicted in Figure 3, where the user request is "Open the Discord application and send Hi to Dylan Li".
-Planner firstly delineates five sub-tasks: "Open Discord", "Navigate to friends", "Search for Dylan Li", "Open Chat with Dylan Li", and "Send a Message", all of which are assigned to Programmer.
-Upon completion of the first subtask by the Programmer, Mentor observes that Discord has been successfully launched and notes that "Dylan is already present in the application." Planner, upon receiving this feedback, revises the initial plan to "Open chat with Dylan" and "Send the message", and then reassigns it to Programmer.
-In this instance, Planner overlooks the fact that programmer needs to specifically integrate Discord’s proprietary API to accomplish these tasks.
-Adapting to the proprietary interfaces of thousands of applications to achieve universality is clearly impractical.
-Furthermore, unforeseen scenarios such as pop-up notifications or advertisements may disrupt task execution.
-Programmer fails to execute and returns the message, "Sorry, I don’t know how to do this." Upon receiving this message, Planner reallocates the task to Viewer.
-Viewer decomposes the coarse-grained task into finer-grained tasks: "click on Dylan", "type Hi", and "return the message", which are then successfully executed.
-Additionally, each subtask completion is reviewed by Mentor, and upon approval, the process advances to the next round.
-The process concludes after all rounds are completed.
-We assess MMAC-Copilot using two benchmarks: GAIA [26] and our newly introduced Visual Interaction Benchmark (VIBench).
-GAIA is designed to test the abilities of general artificial intelligence assistants.
-The MMAC-Copilot performed in level-1, level-2, and level-3 tasks with scores of 45.16, 20.75, and 6.12, respectively.
-It outperforms previous methods by 6.8%.
-VIBench, specifically developed for this study, tests the framework’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office environments.
-The results exhibits outstanding adaptability, handling interfaces with a high degree of accuracy, enhancing interaction within digitally environments.
+Large Language Models (LLMs) Achiam et al.
+(2023); Anil et al.
+(2023); Anthropic (2023); Driess et al.
+(2023); Ouyang et al.
+(2022); Touvron et al.
+(2023a, b) has demonstrated unprecedented capabilities to tackle complex tasks that require human-like reasoning, decision-making, and collaborative efforts Ding et al.
+(2023); Hao et al.
+(2023); Park et al.
+(2023); Qian et al.
+(2023); Xi et al.
+(2023).
+Soon afterwards, Large Vision Models (LVMs) Cai et al.
+(2023); Li et al.
+(2024); Liu et al.
+(2023, 2024); Luo et al.
+(2023); Zhou et al.
+(2023a) expand beyond traditional text-based processing by integrating a visual dimension into LLMs.
+The applications of LVMs have reached diverse domains, with one notable example being their pivotal usage in autonomous agents Wu et al.
+(2024).
+Applying large language models (LLMs) in operating systems faces two primary challenges.
+First, their ability to acquire real-world environmental information is constrained by reliance on a single text modality, limiting the breadth of applications they can effectively manage.
+Second, hallucinations stemming from knowledge domain gaps can lead to flawed planning by LLM agents, with potentially severe consequences.
+Agents typically require both a unified interaction interface and a high level of generalization to adapt to the diverse and complex range of applications within an operating system.
+While previous works Tan et al.
+(2024); Wu et al.
+(2024); Zhang et al.
+(2024) have leveraged scripting and LLMs’ planning abilities to achieve success in certain scenarios, these challenges persist and must be addressed for broader applicability.
+To address the aforementioned challenges, we present MMAC-Copilot, a framework designed to enhance information exchange and collaboration among multi-modal agents.
+Central to this framework is the team collaboration chain, where agents—each driven by an LLM—adapt the initial plan based on their specific domain expertise.
+When a user query is received, the Planner formulates an initial plan based on the query’s requirements.
+The plan is then iteratively refined through collaboration with the other agents, ensuring that insights from each agent’s domain expertise continuously adapt and improve the execution process.
+As illustrated in Figure 1, MMAC-Copilot integrates five key agents:
+  • Planner: Specializes in strategic planning, resource allocation, and ensuring that all agents’ efforts are aligned toward achieving the target.
+  • Librarian: Excels in question-answering tasks and API-based information retrieval.
+  • Programmer: Handles coding tasks, with expertise in executing Bash commands and Python scripts.
+  • Viewer: Focuses on processing visual information, understanding image content, and performing click-based interactions.
+  • Video Analyst: Analyzes video content to extract key events.
+In addition to these core agents, an auxiliary agent, the Mentor, supervises the system’s interactions with the environment, reviewing the user interface after each interaction to confirm whether the desired outcomes have been achieved.
+We evaluate MMAC-Copilot using two benchmarks: GAIA Mialon et al.
+(2023) and our newly introduced Visual Interaction Benchmark (VIBench).
+GAIA is designed to assess general artificial intelligence assistants, and MMAC-Copilot achieved scores of 45.16, 20.75, and 6.12 on level-1, level-2, and level-3 tasks (as detailed in Appendix C), outperforming previous methods by 6.8%.
+VIBench, developed specifically for this study, evaluates performance in non-API-interactable applications across domains such as 3D gaming, recreation, and office environments.
+The results demonstrate MMAC-Copilot’s exceptional adaptability and precision in handling complex interfaces.
 In summary, our main contributions are as follow:
-  • We propose MMAC-Copilot, a collaborative framework involving specialized agents designed to navigate and perform general operating system tasks.
-MMAC-Copilot enhances the interaction capabilities of autonomous virtual agents with operating systems by leveraging multi-modality in processing to tasks.
-The framework consists of distinct modal agents: "Librarian" for question-answering and information retrieval, "Programmer" for executing code and bash commands, "Viewer" for interpreting visual information, and "Video Analyst" for analyzing video content.
-By leveraging the collaboration of modal agents, MMAC-copilot can handle a diverse range of tasks with enhanced accuracy.
-  • To mitigate the incorrect planning issue caused by LLMs’ hallucination, we introduce the team collaboration chain which allows participating agents to adjust the initial plan based on their domain expertise into MMAC-Copilot.
+First, we propose MMAC-Copilot, a collaborative framework involving specialized agents designed to navigate and perform general application tasks.
+MMAC-Copilot enhances autonomous agents’ ability to interact with user interfaces when processing tasks.
+Next, to mitigate the incorrect planning issue caused by LLMs’ hallucination, we introduce the team collaboration chain which allows participating agents to adjust the initial plan based on their domain expertise into MMAC-Copilot.
 It has demonstrated adaptability in handling a variety of case.
-  • We develop the Visual Interaction Benchmark (VIBench), specifically designed to assess the system’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office scenarios.
-VIBench complements existing benchmarks by focusing on the agents’ ability to interact with dynamic and visually complex interfaces, thus providing a more strict test of their practical interaction capabilities.
+Finally, we develop the Visual Interaction Benchmark (VIBench), specifically designed to assess the system’s performance in non-API-interactable applications across diverse domains such as 3D gaming, recreation, and office scenarios.
+VIBench complements existing benchmarks by focusing on the agents’ ability to interact with visually complex user interfaces.
 2 Related work
-In this section, we review some related works, with a focus on the domains of LLM-based agents and Autonomous virtual agents.
-Figure 2: An overview of MMAC-Copilot, below is the concrete definition of each agent.
+In this section, we review some related works, with a focus on the domains of LLM-based agents and autonomous agents.
+Figure 1: An overview of MMAC-Copilot framework.
 2.1 LLM-based agents
-The sunrise field of LLM-based agents, highlighted by projects such as AutoGPT [33] and TaskWeaver [30], marks a significant milestone in the field of LLM research.
+The sunrise field of LLM-based agents, highlighted by projects such as AutoGPT Significant-Gravitas (2023) and TaskWeaver Qiao et al.
+(2023), marks a significant milestone in the field of LLM research.
 These agents adopt a human-like approach to problem-solving by decomposing complex tasks into manageable subtasks.
-The capability demonstrates a profound advancement in planning, observation, and responsive actions, closely mirroring human cognitive processes [38, 40, 45].
-The LangChain Agent [36] further extends these capabilities by enabling LLMs to sequence actions, thus expanding the potential for automation across various domains such as customer service, logistics, and content creation.
+The capability demonstrates a profound advancement in planning, observation, and responsive actions, closely mirroring human cognitive processes Touvron et al.
+(2023b); Wang et al.
+(2024); Xi et al.
+(2023).
+The LangChain Agent Topsakal and Akinci (2023) further extends these capabilities by enabling LLMs to sequence actions, thus expanding the potential for automation.
 This strategic utilization of LLMs underscores a move towards more sophisticated, context-aware AI systems.
-Developments in multi-agent systems, such as AutoGen [43], MetaGPT [17], and AutoAgents [8], introduce a novel dimension to the field by enabling collaboration and competition among agents.
+Developments in multi-agent systems, such as AutoGen Wu et al.
+(2023), MetaGPT Hong et al.
+(2023a), and AutoAgents Chen et al.
+(2023), introduce a novel dimension to the field by enabling collaboration and competition among agents.
 This architecture leverages the unique strengths of individual agents, enhancing system-wide efficiency and adaptability in complex scenarios.
-Such collaborative frameworks are proving essential for large-scale applications in sectors like healthcare [10, 20], finance [14], and urban planning [4].
-The versatility and efficacy of these LLM-based agents are highlighted in a range of application domains: In robotics, agents assist in both navigational tasks and complex manipulations, contributing to more autonomous and adaptive robotic systems [12].
-Web manipulation and data extraction have seen significant enhancements, enabling better data-driven decision-making and user interaction models [49, 55].
-The gaming industry benefits from more dynamic AI opponents and game management systems, which improve user engagement and game design [13, 35].
-Automated data analysis tools powered by LLM agents offer sophisticated insights and predictions, facilitating more effective data handling in business and research [39, 53].
-2.2 Autonomous virtual agents
+Such collaborative frameworks are proving essential for large-scale applications in sectors like healthcare Dhatterwal et al.
+(2023); Jones et al.
+(2022), finance Gautier et al.
+(2023), and urban planning Bae et al.
+(2022).
+The versatility and efficacy of these LLM-based agents are highlighted in a range of application domains: In robotics, agents assist in both navigational tasks and complex manipulations, contributing to more autonomous and adaptive robotic systems Driess et al.
+(2023).
+Web manipulation and data extraction have seen significant enhancements, enabling better data-driven decision-making and user interaction models Yao et al.
+(2022); Zhou et al.
+(2023b).
+The gaming industry benefits from more dynamic AI opponents and game management systems, which improve user engagement and game design Fan et al.
+(2022); Tan et al.
+(2024).
+Automated data analysis tools powered by LLM agents offer sophisticated insights and predictions, facilitating more effective data handling in business and research Wang et al.
+(2023); Zhang et al.
+(2023b).
+2.2 Autonomous agents
 The application of LVM systems for interfacing with Graphical User Interfaces (GUIs) in digital applications have been a key area of research.
-Wonderland [46] harnessed the capabilities of GPT-4V which delineated by Yang et al.
-[47] to navigate through mobile application interfaces by analyzing screenshots.
-Concurrently, the AppAgent [48] utilizes GPT-4V to simulate the actions of mobile users, facilitation the execution of tasks in mobile applications through the analysis of device snapshots.
+Wonderland Yan et al.
+(2023) harnessed the capabilities of GPT-4V which delineated by Yang et al.
+Yang et al.
+(2023b) to navigate through mobile application interfaces by analyzing screenshots.
+Concurrently, the AppAgent Yang et al.
+(2023a) utilizes GPT-4V to simulate the actions of mobile users, facilitation the execution of tasks in mobile applications through the analysis of device snapshots.
 Moreover, the MobileAgent incorporates Optical Character Recognition (OCR) technology to enhance the capabilities of of GPT-4V within a similar mobile agent architecture.
 This strategic enhancement allows MobileAgent to attain task completion rates on par with human operators.
-In contrast, CogAgent [18] constructs a specifically engineered LVM for interpreting and navigating GUIs.
-UFO [51] utilizes the windows inspect tool to acquire component and control information from system applications, providing it as context to GPT-4V to fully leverage its planning capabilities.
+In contrast, CogAgent Hong et al.
+(2023b) constructs a specifically engineered LVM for interpreting and navigating GUIs.
+UFO Zhang et al.
+(2024) utilizes the windows inspect tool to acquire component and control information from system applications, providing it as context to GPT-4V to fully leverage its planning capabilities.
 3 MMAC-Copilot Framework
@@
 11:         Go to step 4 with revised $P_{i}$
-12:     end if
-13:end for
-14:Verify if $S_{i}$ satisfies $R$
-15:if not satisfied then
-16:     Retry
-17:else
-18:     Break loop and finalize task
-19:end if
+12:     else if $F_{i}$ indicates no adjustments are needed then
+13:         if $s$ is the last subtask in $P_{0}$ then
+14:              Terminate loop
+15:         else
+16:              Continue to the next subtask
+17:         end if
+18:     end if
+19:end for
 3.1 Specialization of agents
-Empirical studies suggest that diversity within human groups contributes to a variety of perspectives, which in turn enhances the group’s performance across various tasks [6, 41, 42].
-Also, contemporary research indicates that assigning specific roles to autonomous agents can improve their effectiveness [22, 29, 31].
+Empirical studies suggest that diversity within human groups contributes to a variety of perspectives, which in turn enhances the group’s performance across various tasks Bransford and Stein (1993); Williams and O’Reilly III (1998); Woolley et al.
+(2015).
+Also, contemporary research indicates that assigning specific roles to autonomous agents can improve their effectiveness Li et al.
+(2023); Qian et al.
+(2023); Salewski et al.
+(2024).
 Thus, in our framework, we have defined six agents, namely: Planner, Librarian, Programmer, Viewer, Video Analyst, and Mentor, as shown in Figure 1.
 The followings are detailed introductions to each agent:
-  • Planner: Employs GPT-4 to strategically manage and allocate tasks among agents, optimizing workflow efficiency.
+Planner: Employs LLM to strategically manage and allocate tasks among agents, optimizing workflow efficiency.
 In contrast to other frameworks, our Planner is not required to decompose tasks into indivisible atomic subtasks.
@@
 It allows each agent to leverage its own expertise, enhancing the adaptability of the overall framework.
-  • Librarian: Utilizes GPT-4 and API for information retrieval, enabling it to answer queries and provide foundational knowledge.
+Librarian: Utilizes LLM and API for information retrieval, enabling it to answer queries and provide foundational knowledge.
 It possesses the capability to rapidly process vast amounts of information, thereby supporting user requests and providing related context with other agent.
-  • Programmer: Responsible for coding and executing scripts, including Python and Bash commands, to interact with software environments directly.
-The Programmer’s role is crucial in implementing precise operating systems interactions that adapt to diverse requirements.
-Key to the Programmer’s efficacy is its ability to refine existing code through a mechanism designed to diagnose and resolve issues effectively.
-When code errors are identified, Programmer analyzes the source of failure, applies necessary modifications, and maintains the original code structure as much as possible.
-This refined process includes a detailed explanation of identified issues, enhancing the clarity and functionality of the code without introducing new exception, and improving the handling of existing ones.
-Furthermore, Programmer can generate accurate Python invocation statements.
-By deeply understanding the context of the provided Python class, analyzing task descriptions, and interpreting method parameter details to ensure the correct application.
-The result is a syntactically correct Python statement that calls the methods with appropriate parameters.
-Lastly, Programmer’s capabilities are complemented by a critical evaluation mechanism.
-This mechanism assesses the refined and executed code against the user’s specified task requirements.
-It evaluates the code’s functionality, alignment with the task, and feedback provided by the environment, culminating in a comprehensive reasoning process.
-The interactions within the Programmer’s mechanism can be mathematically represented as follows:
+Programmer: Be responsible for writing and executing scripts, such as Python and Bash, to directly interact with software environments.
+This role is essential for tailoring applications to meet diverse requirements.
+The Programmer refines existing code by diagnosing and resolving issues.
+This process involves clarifying identified problems, enhancing code functionality, handling exceptions better, and providing detailed explanations.
+Accurate Python statements are generated by understanding the context of Python classes and analyzing method parameters, ensuring syntactically correct calls.
+Also a evaluation mechanism further complements the Programmer’s capabilities, assessing the refined code against task requirements, its functionality, and alignment, leading to a comprehensive review process.
+More detailed Prompt has been listed on Appendix B.
+The interactions within the Programmer’s mechanism are mathematically represented as:
 […]
@@
   • $C_{\text{init}}$ is the initial code.
-      $E$ represents the error analysis and feedback.
-      $R$ is refinement mechanism.
-      $C_{\text{mod}}$ is the modified code.
-      $\delta$ includes environmental variables such as the current working directory or system configurations.
-      $X$ represents execution function.
-      Out includes the runtime output and any runtime errors.
-      $T$ is the original task requirements.
-      $V$ represents the evaluation mechanism.
-      Judge is a boolean indicating if the task is accomplished.
-      Score is an integer rating the code’s generality and robustness. If the score is above the specified threshold, the code for this run is retained.
-  • Viewer: Integrates GPT-4V for interpreting complex visual information from screenshots.
-Coupled with SeeClick [9] model, Viewer could translates visual analysis into actionable commands, such as click on identified UI elements, bridging the gap between visual understanding and operating systems interaction.
-  • Video Analyst: Uses Gemini Vision to process and analyze video content, extracting critical visual data.
-It is significant for subtasks which depend on video-based information, enhancing the model’s ability to generalize across different modal.
-Video Analyst interprets queries, retrieves relevant details from video streams, and delivers accurate responses.
-These capabilities are crucial for providing contextual insights, assisting other agents in understanding the current statement.
-In summary, Video Analyst is adept at handling complex scenarios where video data is key.
-  • Mentor: Powered by GPT-4V, Mentor provides strategic oversight and troubleshooting support after the execution processes.
-To be specific, Mentor firstly analyzes screenshots of the current application window to provide detailed description of visible state.
-And then Mentor identified whether the intended change or commands have successfully impacted the application as expected.
-Additionally, Mentor generates insights which are derived from a deep analysis of the outcomes and are structured as feedback to be incorporated into the next round.
+  • $E$ stands for error analysis.
+  • $R$ is the refinement process.
+  • $C_{\text{mod}}$ represents the modified code.
+  • $\delta$ includes environmental variables.
+  • $X$ is the execution function.
+  • Out includes runtime outputs and errors.
+  • $T$ corresponds to the original task requirements.
+  • $V$ is the evaluation process.
+  • Judge determines if the task is accomplished.
+  • Score assesses the code’s generality and robustness, with higher scores indicating retention.
+Viewer: The Viewer module integrates GPT-4V for the sophisticated interpretation of complex visual data extracted from screenshots.
+By coupling this capability with the SeeClick model Cheng et al.
+(2024), the Viewer effectively translates visual analyses into executable commands—such as initiating clicks on identified user interface elements.
+This integration bridges the gap between visual comprehension and application interaction, enabling automated manipulation of graphical interfaces based on visual understanding.
+Video Analyst: The Video Analyst leverages the Gemini Vision to process and analyze video content, extracting critical visual information relevant to the task domain.
+This module is essential for subtasks that depend on video-based data, enhancing the model’s ability to generalize across different modalities.
+By interpreting queries and retrieving pertinent details from video streams, the Video Analyst delivers precise and contextually appropriate responses.
+These capabilities are crucial for providing comprehensive contextual insights and assisting other agents in understanding the current operational state.
+Consequently, the Video Analyst adeptly handles complex scenarios where video data is paramount.
+Mentor: Powered by GPT-4V, the Mentor agent plays a pivotal role in overseeing the system’s execution processes, providing strategic oversight and troubleshooting support.
+After each subtask is executed, the Mentor analyzes screenshots from the current application window, offering a detailed description of the visible state.
+It evaluates whether the intended changes or commands have successfully impacted the application as expected.
+If the analysis indicates that adjustments are needed, the Mentor generates feedback based on these observations, which is then incorporated into the system’s planning process for the next iteration.
+This feedback loop allows the planner to revise the current plan and reattempt the adjusted task.
+However, if no adjustments are necessary, the system proceeds to the next subtask in the sequence.
+In the case where the last subtask is reached and no further changes are required, the system terminates the process.
+This iterative mechanism of feedback and improvement not only ensures the successful completion of tasks but also enhances the overall system performance by continuously refining actions based on empirical data.
 3.2 Communication Protocols
-Following established multi-agent frameworks [17, 22, 28, 52, 56], MMAC-Copilot adopts Structured Communication Interfaces, avoiding the traps of unconstrained nature language communication.
+Following established multi-agent frameworks Hong et al.
+(2023a); Li et al.
+(2023); Park et al.
+(2023); Zhang et al.
+(2023a); Zhuge et al.
+(2023), MMAC-Copilot adopts Structured Communication Interfaces, avoiding the traps of unconstrained nature language communication.
 Each agent communicates through a defined schema appropriate for their role.
@@
 For example, "observation" key is exclusive output by Mentor and Viewer, as it is specifically to their role in emphasizing visual observations of screenshots.
-Apart from themselves, only the Planner receives the content of "observation" key, both when formulating initial plans and when revising them.
+Apart from themselves, only the Planner receives the content of "observation" key, both when revising initial plans.
 These protocols are designed to prevent common issues such as data ambiguity in complex multi-agent interactions.
@@
 By distributing responsibilities and decision-making power, these chains enable a more agile and responsive organizational environment.
-Such chain has been successfully implemented across multiple domains [5, 19, 32, 50].
+Such chain has been successfully implemented across multiple domains Bang and Dalsgaard (2005); Huh et al.
+(2019); Selvanathan and Poravi (2018); Yatsuka et al.
+(2020).
 In the MMAC-Copilot framework, team collaboration chain is essential for managing difficult tasks.
@@
 3.3.1 Dynamic Planning and Subtask Refinement
-During the planning phase, Planner uses input from the Viewer to formulate an initial plan.
+During the planning phase,In the planning phase, the planner formulates an initial plan based on user requests.
 Initial plan often outline the subtasks in a coarse manner, addressing the dynamic nature of visual environments and the inherent limitations of planning based on textual information alone.
@@
 This feedback is crucial for revising the current plans and is incorporated into the next round of planning.
-This feedback loop between the Mentor and the Planner not only ensures high-quality task completion but also facilitates adaptive learning, allowing the MMAC-Copilot system to evolve its strategies based on real-world interactions and outcomes.
 4 Experiments
@@
 4.1 Benchmark and setting
-We evaluate the performance of MMAC-Copilot using two distinct benchmarks designed to challenge and assess the capabilities of AI assistants in diverse interaction environments.
-The first is the General AI Assistant Benchmark (GAIA) [26], which includes 466 rigorous question-answering tasks.
+Table 1: Performance comparison on GAIA. All performance are reported on the private test set.
+Model Level 1 Level 2 Level 3 Average
+Human* 93.90 91.80 87.30 91.00
+Search engine 7.4 0.00 0.00 2.47
+GPT-3.5 Achiam et al. (2023) 4.3 1.89 2.08 2.67
+GPT-4 Achiam et al. (2023) 9.68 1.89 0.00 4.00
+AutoGPT4 Significant-Gravitas (2023) 15.05 0.63 0.00 5.00
+GPT4 Turbo Achiam et al. (2023) 9.68 6.92 0.00 6.67
+GPT-4 Plugins Achiam et al. (2023) 30.30 9.70 0.00 14.60
+Chamomile Guillot (2023) 16.13 17.61 2.08 14.67
+FRIDAY Wu et al. (2024) 40.86 20.13 6.12 24.25
+Ours 45.16 20.75 6.12 25.91 (+6.8%)
+Table 2: Evaluation results on VIBench.
+Model 3D Recreation Office Average
+UFO 0.00 28.57 15.38 14.65
+FRIDAY 31.58 42.86 30.77 35.07
+Ours 63.16 69.23 78.57 70.32 (+35.25)
+We evaluate the performance of MMAC-Copilot using two distinct benchmarks designed to challenge and assess the capabilities of autonomous agent in diverse interaction environments.
+The first is the General AI Assistant Benchmark (GAIA) Mialon et al.
+(2023), which includes 466 rigorous question-answering tasks.
 GAIA primarily tests AI assistants in a question-answering (QA) format, focusing on their ability to utilize API calls for information retrieval and task execution.
@@
 Unlike GAIA, VIBench emphasizes the necessity for agents to interact with applications through graphical user interfaces (GUIs) that cannot typically be controlled via standard APIs such as Win32 APIs.
-Mathematical Representation of VIBench
 The core interaction process in VIBench can be mathematically described as follows: Let $S_{0}$ represent the initial system state which includes the current GUI appearance and available actions.
@@
 Objective:Verify that $\text{Objective}:\text{Verify that }S_{f}\text{ meets }R$
-VIBench thus complements GAIA by adding a layer of complexity and realism, challenging agents to not only respond correctly in a QA format but also to effectively navigate and interact within software environments that mirror everyday user challenges.Examples are shown in Figure 4
+VIBench thus complements GAIA by adding a layer of complexity and realism, challenging agents to not only respond correctly in a QA format but also to effectively navigate and interact within a application environments that mirror everyday user challenges.
+Examples are shown in Figure 3.
+We also present the detailed list of request utilized in VIBench in Appendix D.1.
 Setting
@@
 For the Video Analyst, we used gemini-1.5-pro-vision.
+More detailed experiment setting could been seen in Appendix C.
 4.2 Evaluation Method
 For GAIA, each question calls for a specific response: string, number, or floats. Evaluation is conducted through a exact match between the model’s answer and ground truth.
-For our VIBench, we adopt human evaluation which follow the SIMA evaluation method [34], recognizing that in numerous instances, task success cannot be automatically inferred.
+For our VIBench, we adopt human evaluation which follow the SIMA evaluation method SIMA Team (2024), recognizing that in numerous instances, task success cannot be automatically inferred.
 We engage human judges who possess expert-level familiarity with the specific applications, defined as having at least 20 hours of application usage.
@@
 The selection of Plugins for GPT-4 is conducted manually, with individual choosing based on specific user query.
-For comparative purposes, we also include data on human performance sourced from GAIA [26].
+For comparative purposes, we also include data on human performance sourced from GAIA Mialon et al.
+(2023).
 For VIBench, we report the performance of UFO and FRIDAY.
@@
 4.4 Results
-The results of our experiments with MMAC-Copilot provide a comprehensive assessment of its capabilities in handling complex, multi-modal tasks with in digital environments.
-In this session, we present and discuss the findings from our evaluations on both the GAIA benchmark and our newly developed Visual Interaction Benchmark(VIBench).
-Table 1: Performance comparison on GAIA. All performance are reported on the private test set.
-Model Level 1 Level 2 Level 3 Average
-Human* 93.90 91.80 87.30 91.00
-GPT-3.5 [1] 4.3 1.89 2.08 2.67
-GPT-4 [1] 9.68 1.89 0.00 4.00
-AutoGPT4 [33] 15.05 0.63 0.00 5.00
-GPT4 Turbo [1] 9.68 6.92 0.00 6.67
-GPT-4 Plugins [1] 30.30 9.70 0.00 14.60
-Chamomile [15] 16.13 17.61 2.08 14.67
-FRIDAY [44] 40.86 20.13 6.12 24.25
-Ours 45.16 20.75 6.12 25.91 (+6.8%)
 MMAC-Copilot shows significant performance improvement across all levels of GAIA.
@@
 Given that tasks in GAIA encompass various complex AI assistant capabilities, from basic information retrieval to interactive task requiring multiple steps and integration of external data sources.
-Table 2: Evaluation results on VIBench.
-Model 3D Gaming Recreation Office Average
-UFO [15] 0.00 28.57 15.38 14.65
-FRIDAY [44] 31.58 42.86 30.77 35.07
-Ours 63.16 69.23 78.57 70.32 (+35.25)
-In VIBench as shown in Table LABEL:tab:VIBench, MMAC-Copilot outperformed other systems across all application categories (3D gaming, recreation, and office scenarios).
-Our model achieved average scores fo 63.16% in 3D gaming, 69.23% in recreation, and 78.57% in office applications, with an overall average of 70.32%.
+In VIBench as shown in Table 2, MMAC-Copilot outperformed other systems across all application categories (3D gaming, recreation, and office scenarios).
+Our model achieved average scores of 63.16% in 3D gaming, 69.23% in recreation, and 78.57% in office applications, with an overall average of 70.32%.
 It indicates MMAC-Copilot’s robust capability in navigating and interacting within complex GUI environments, where direct API interaction is not possible.
-MMAC-Copilot’s superior performance can be attributed to its effective utilization of image and video analysis, enabling precise navigation and interaction through the GUIs of different software.
 5 Conclusion
-This paper presents MMAC-Copilot, a novel framework designed to enhance the interaction capabilities of autonomous virtual agents with operating systems.
-Our approach capitalizes of autonomous virtual agents with operating systems.
-Our approach capitalizes on a team collaboration chain that harnesses the collective expertise of each agents, including "Planner", "Librarian", "Programmer", "Viewer", "Video Analyst", and "Mentor".
-Each agent brings a unique set of skills and perspectives, enabling an adaptive interaction with complex digital environments.
-Also MMAC-Copilot innovatively mitigates the issues with hallucination among the agents.
-By facilitating dynamic and team collaboration among the agents, our framework allows for continuous adaptation and refinement of plans based on environmental changes.
-The performance of MMAC-Copilot has been tested on GAIA benchmark and our newly introduced VIBench, showing significant improvements over existing state-of-the-art systems.
-6 Discussion
+MMAC-Copilot is a novel framework enhances agent interactions within complex digital environments.
+This framework effectively addresses agent hallucination, ensuring reliable responses to environmental changes.
+Tested on the GAIA benchmark and our VIBench, MMAC-Copilot demonstrates superior performance compared to existing systems.
+Limitation
 We acknowledge that while MMAC-Copilot has achieved a level of generality in application through the integration of GUI models for visual detection, it is still limited by the GUI model’s ability to understand UI components.
 Consequently, MMAC-Copilot struggles with interpreting complex UI interfaces and performing multi-step operations that depend on such complexity.
-Additionally, MMAC-Copilot’s capabilities in real-time 3D gaming environments reveal gaps in the spatial understanding, which is cruical for tasks such as navigation, aiming and shooting in 3D environments.
+Additionally, MMAC-Copilot’s capabilities in real-time 3D gaming environments reveal gaps in the spatial understanding, which is crucial for tasks such as navigation, aiming and shooting in 3D environments.
 The framework’s longer inference time impair the ability to perform tasks with higher temporal sensitivity.
@@
 This approach will not only help in understanding complex UI layouts but also in recognizing user intentions and environmental cues in real-time games.
+Appendix A Demo on Mac OS
+The following Figure 4 illustrates an example of MMAC-Copilot performing an action sequence on macOS in response to the user request “Play <Love Story> on Spotify.” Each subfigure represents a distinct step in the process: (A) shows the “Planner” formulating the initial plan; (B) the “Programmer” launching the application using a bash command; (C) the “Viewer” clicking the search box and entering <Love Story>; and (D) the “Viewer” clicking the play button.
+This demonstrates how MMAC-Copilot efficiently handles a media playback request from the user.
+Appendix B Agent prompts
+To better explain the operating mechanism of Programmer, we have provided the complete prompts in the Figure 9,10,11,12,13,15,16,17 and 18.
+Appendix C GAIA Detail Experiment Setting
+GAIA serves as a comprehensive benchmark for evaluating general AI assistants by testing their ability to handle real-world scenarios across various capabilities, including advanced reasoning, multi-modality handling, coding, and web browsing.
+The benchmark comprises text-based questions, sometimes accompanied by supplementary files such as images or spreadsheets.
+These questions are designed to cover a broad range of use cases from everyday tasks to complex scientific queries, necessitating concise and unambiguously correct answers that are straightforward to verify.
+C.1 Levels of Difficulty
+The questions in GAIA are categorized into three levels based on their complexity:
+  • Level 1: Simplest tasks, generally requiring minimal tool use and up to five steps.
+  • Level 2: More complex tasks that require integrating information from multiple sources, involving about 5 to 10 steps.
+  • Level 3: The most complex tasks, demanding extensive sequences of actions and the use of multiple tools to solve real-world problems.
+We present a sample GAIA question from Level 1 through Level 3 in Figure 5
+Figure 5: Sample Questions from GAIA.
+These tasks necessitate core competencies including reasoning, the ability to handle multiple modalities, and proficiency in using tools.
+The answers to these questions are unequivocally defined and are intentionally unlikely to be present in unstructured training datasets.
+Several questions are supplemented with visual evidence, mirroring authentic scenarios and enhancing the precision of query formulation.
+C.2 Evaluation Methodology
+GAIA employs an automated, precise, and factual evaluation method.
+Each question demands an answer in a specified format (string, number, or comma-separated list) and correctness is verified through an exact match with a pre-defined ground truth, ensuring the evaluation’s reliability and efficiency.
+C.3 Model Experimental Setup
+In this section, we introduce the detailed experiment setting of our comparison models.
+We evaluates leading AI models, including GPT-3.5, GPT4 and AutoGPT, also conduct non-AI baselines experiments such as human annotators and traditional web search methods.
+We will introduce each experiment setting as following.
+Comparison Models:
+GPT3.5/GPT4:
+This setup evaluates the core linguistic and reasoning capabilities of GPT4 without the aid of any external tools or plugins.
+We use a prefix prompt before asking the model a question.
+To ease answer extraction, we specify a format in the prefix prompt which could been saw in Figure 6.
+GPT4 with Plugins:
+In this configuration, GPT4 is enhanced with selected plugins that extend its functionality. This includes tools for advanced data analysis, web browsing, and handling of third-party APIs.
+AutoGPT:
+Utilizing GPT4 as its backend, AutoGPT automatically selects plugins based on the context of the query.
+This model is tested for its efficiency in tool selection and its ability to autonomously adapt to complex query requirements.
+Human*:
+Consisting of experts in various fields, human annotators provide responses based on their expertise and research abilities.
+Their performance sets a high standard for AI models, especially in tasks requiring deep domain knowledge and nuanced understanding.
+Search engine:
+Using conventional search engines, this baseline evaluates the effectiveness of straightforward information retrieval.
+Specifically, we enter our questions into a search engine and review the first page of results to see if the answer can be inferred.
+Notes: Each model undergoes rigorous testing across a series of structured tasks that range from simple to highly complex, categorized into three levels of difficulty as previously described.
+Models are evaluated based on accuracy, the relevance of the responses, and the efficiency of handling the queries.
+Each session is replicated three times to ensure consistency and reliability in the results.
+Figure 6: The system prompt to answer GAIA.
+Appendix D VIBench
+D.1 VIBench Results List
+In Table 3, we provide a comprehensive list of the user requests featured in VIBench, along with the corresponding performance of MMAC-Copilot.
+These requests cover a range of commonly used functions across eight popular applications.
+For clarity, requests that involve follow-up tasks are numbered sequentially.
+In the Success column, “” denotes that MMAC-Copilot completes the request successfully, while “” indicates a failure to fufill the request.
+Table 3: The requests in VIBench and detailed results achieved by MMAC-Copilot
+Operation Application Success
+Ride on vehicle [Press F] PUBG
+Probe shot [Press Left mouse] PUBG
+Picking up supplies [Identify the screen and press F] PUBG
+Assemble accessories and discard unnecessary ones PUBG
+Use E skill Valorant
+Shoot Valorant
+Climb up the rope Valorant
+Absorb energy ball Valorant
+Take turns to release Q skills Genshin
+Pick up the Divine Eye Genshin
+Freeze trap Genshin
+Fishing Genshin
+Climb Genshin
+Send message to xxx DingTalk
+Create meeting DingTalk
+Create a group DingTalk
+End meeting TencentMeeting
+Turn on camera TencentMeeting
+Create a meeting TencentMeeting
+Join a meeting TencentMeeting
+Click the record bottom TencentMeeting
+Copy Invitation TencentMeeting
+Open the library in steam Steam
+Open the friends interface Steam
+Click on store and then click on the search box Steam
+View current music and play it Spotify
+Search for a specific piece of music and save it Spotify
+Open docuseries in Netflix Netflix
+Save XXX Series to my list Netflix
+Remove XXX from my list Netflix
+D.2 Visualization
+We have visualized VIBench with requests in Figure 7 and 8.
+D.3 Adapt for Windows
+To compatible with the Windows platform, we made some adaptations focusing on tool integration and modifications for platform-specific libraries.
+One of the primary adaptations involved the integration of tools such as the Python environment and Windows Shell support.
+The Python interpreter was configured to work within the Windows platform and Windows path handling were incorporated to ensure smooth execution of scripts and Python modules.
+In addition to tool integration, some Python libraries required adaptation to function properly on Windows.
+We used alternative Python modules to replace database libraries that did not fully support the Windows environment for tool persistent.
+As Python libraries behave differently on Windows due to variations in concurrent requests management, we also made some adaptation for it.
+Figure 7: How VIBench works(Part-I)
+Figure 8: How VIBench works(Part-II)
+Figure 9: Prompt for Programmer Part
+Figure 10: Prompt for Programmer Part
+Figure 11: Prompt for Programmer Part
+Figure 12: Prompt for Programmer Part
+Figure 13: Prompt for Programmer Part
+Figure 14: Prompt for Programmer Part
+Figure 15: Prompt for Programmer Part
+Figure 16: Prompt for Programmer Part
+Figure 17: Prompt for Programmer Part
+Figure 18: Prompt for Programmer Part